OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $14,000. Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $2,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.)

jimthompson5910 (jim_thompson5910):

option 1 divide the total cost of 14,000 by 60 months to find the monthly payment

jimthompson5910 (jim_thompson5910):

oh wait, for option 1, you made a down payment of $2,000

jimthompson5910 (jim_thompson5910):

so you are dividing 12,000 by 60

jimthompson5910 (jim_thompson5910):

for option 2, you first calculate the interest to get I = Prt I = 14000*0.02*6 I =1680 Then you add this onto the amount borrowed to get 14000+1680= 15680 Afterwards, you divide 15680 by 60 to get the monthly payment for option 2